rebeccaxhawaii (rebeccaxhawaii):

"Completing the square involves dividing the coefficient on x by 2 and squaring the result" but why

OpenStudy (welshfella):

one reason we complete the square is to solve a quadratic equation that cannot be factored. An example is as follows solve for x x^2 - 2x - 7 = 0 x^2 - 2x = 7 divide -2 by 2 and square the results (x - 1)^2 - 1 = 7 we needed to subtract the 1 to make the left side equal to x^2 - 2x (x - 1)^2 = 8 now find the square root of both sides to solve for x

OpenStudy (anonymous):

Let's work backwards. Here's one part of our completing the square answer. (x+a)^2 Expand this... (x+a)(x+a) = x^2 + ax + ax + a^2 Simplifies to: = x^2 + 2ax + a^2 See how the coefficient (a) is double the number we had in our answer?